OpenStudy (anonymous):

Angle EFG and angle GFH are a linear pair, mAngle EFG=4n+25, and mAngle GFH=5n+29. What are mAngleEFG and mAngleGFH?

OpenStudy (mertsj):

Do you know what it means for the angles to be a linear pair?

OpenStudy (anonymous):

Doesn't it mean that it's a pair of adjacent angles whose noncommon sides are opposite rays?

OpenStudy (mertsj):

Yes. More importantly, it also means that the sum of the two angles is 180. Does that help you?

OpenStudy (anonymous):

Kinda. But i still don't know how i would solve this problem

OpenStudy (mertsj):

One angle is 4n+25 The second angle is 5n + 29 Can you add those two?

OpenStudy (anonymous):

wouldn't i just add 4 to 25? and get 29?

OpenStudy (mertsj):

4n+25+5n+29=9n+54

OpenStudy (mertsj):

So 9n+54=180

OpenStudy (mertsj):

Solve that equation.

OpenStudy (anonymous):

would i subtract 9 from each side?

OpenStudy (mertsj):

1. subtract 54 from both sides. 2. Divide both sides by 9

OpenStudy (anonymous):

Ohhh okay! one second

OpenStudy (anonymous):

so i got 14

OpenStudy (mertsj):

You got n = 14. The problem asked you to find the measure of each angle.

OpenStudy (anonymous):

Well for the equation 9n+54=180 I got n=14

OpenStudy (mertsj):

What does the problem say about the measure of angle EFG?

OpenStudy (anonymous):

Angle EFG and angle GFH are a linear pair, mAngle EFG=4n+25, and mAngle GFH=5n+29. What are mAngleEFG and mAngleGFH? m<EFG=___ m<GFH=___ that's all the questions says

OpenStudy (mertsj):

Read the problem and answer my question. What does the problem say about the measure of angle EFG?

OpenStudy (mertsj):

mAngle EFG=4n+25

OpenStudy (anonymous):

Yeah was just about to say that

OpenStudy (mertsj):

Then find the measure of EFG. You know what n is.

OpenStudy (anonymous):

ohh okay, i'm getting it once second let me work it out

OpenStudy (anonymous):

81?

OpenStudy (mertsj):

Yes. Now find the measure of the other angle. Remember: the two angles are a linear pair which means their sum is 180.

OpenStudy (anonymous):

99? For the other equation

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

okay so m<EFG=81 m<GFH=99

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Okay, Thank you so much!